# Raman Microspectroscopy for Real-Time Structure Indicator in Ultrafast Laser Writing

### v1

@@ Abstract
Raman microspectroscopy provides non-destructive, structure-specific contrast suited to in situ monitoring of femtosecond (fs) laser fabrication processes. In this work, we introduce a sample-agnostic Raman workflow that screens candidate spectral indicators using drift-resistant, self-normalized metrics and, where appropriate, label-free hyperspectral unmixing to extract endmembers and validate metric selection. Using diamond as an example, hyperspectral maps and electrical tests show that depletion of the 1332 cm-1 sp3 line is a monotonic, robust predictor of resistance—counterintuitively outperforming sp2-based measures often assumed optimal. This enables a practical implementation: a single narrowband channel at the selected marker provides real-time feedback without full-spectrum acquisition. Because the selection hinges on general band definitions and functional validation rather than diamond-specific physics, the approach translates Raman phase information directly into device specifications and is transferable to other Raman-active hosts for in-line quality control of laser-written structures.
I Introduction
Femtosecond laser writing enables direct, three-dimensional patterning of embedded features in transparent hosts via nonlinear absorption—multi-phonon ionization (MPI) and Zener breakdown—that initiates localized structural transformation [1, 2, 3]. These processes permit buried and surface carbon wires and electrode architectures with arbitrary geometry for electrically conductive devices [4], including radiation detectors [5]. Practical subsurface fabrication benefits from pre-compensating high-index–induced spherical aberration using adaptive optics such as a spatial light modulator (SLM) to restore a tight focal volume [6, 7]. While the resulting microstructures have been examined by destructive or ex-situ methods—TEM [8, 9], SEM [2], and X-ray diffraction microscopy (XRD) [10]—their complexity and inability to operate in situ limit process monitoring.
Raman microspectroscopy is a vibrational spectroscopic technique based on the inelastic scattering of monochromatic light. It has been extensively used for non-destructive identification, structural characterization, and monitoring of chemical and physical properties across materials science, chemistry, and biology[11]. In carbon systems in particular, Raman can distinguish bonding configurations, crystallinity, and local strain, making it suitable for non-invasive, in situ, spatially resolved monitoring during fabrication [12].
To obtain robust, comparable indicators from spectra acquired across fields of view and samples, previous studies have adopted a deconvolution and normalization strategy that favors self-referenced, phase-aware metrics over single-peak intensities [11]. Absolute intensities are sensitive to power drift, focus/NA mismatch, baseline fluorescence, and strain; by contrast, combining phase-informative windows and leveraging label-free spectral deconvolution stabilizes the readout and enables cross-instrument comparison. In practice, we explicitly assign bands, normalize to pristine-material pixels, and compute complementary indicators of conversion (ratios and depletion measures defined below), converting hyperspectral maps into quantitative metrics suitable for in situ monitoring [12].
Diamond is an instructive case for Raman-guided fs fabrication. As sp3-bonded carbon, it is an electrical insulator with exceptional hardness, high thermal conductivity, a wide bandgap, and chemical robustness, enabling applications in high-power electronics, photonics, and radiation detection. Point defects such as nitrogen vacancy centers further extend functionality to quantum information and nanoscale sensing [13, 14, 15, 16]. The broader carbon design space—sp2-bonded forms (graphite, graphene, carbon nanotubes) with delocalized electrons for high electrical/thermal conductivity [17, 18], allotropic variants such as hexagonal diamond (lonsdaleite) [19], and mixed-phase nanocomposites such as ’diaphite’ [9, 20, 21]—underscores the value of a phase-resolved optical readout. Although diamond is transparent at 520 nm (2.38 eV) due to its 5.47 eV bandgap, which prevents linear photon absorption from the writing laser, the fabrication process instead relies on nonlinear absorption involving multi-phonon ionization (MPI) and Zener breakdown, initiating a confined sp3 to sp2 phase transition [1, 2, 3], forming local graphitized regions with device-relevant geometries.
@@ I Introduction
Here, we use Raman microspectroscopy to systematically assess fs-laser-written graphitic electrodes in diamond. Leveraging these vibrational modes, we perform direct mapping of surface graphitization through explicit band assignment and label-free spectral deconvolution, and we correlate the spectral signature of the electrodes, specifically the relative intensities of the sp2 band, sp3 line and sp2 to sp3 ratio, with measured electrical conductivity. In doing so, we demonstrate a practical pathway to in-line process monitoring and optimization for laser-written diamond electrodes.
II Results and Discussion
@@
The laser-written graphitic pads and wires were chosen to emulate the two principal conductive elements found in devices. Figure 1a illustrates the experimental concept: a train of fs pulses (400 fs) is scanned through the surface of the diamond while a continuous-wave laser excited the same region and backscattered Raman emission is collected. Pads wider than 25 m and 200 m long reproduce the large‐area contact zones required for wire bonding or subsequent metallization, where sheet resistance and adhesion set the performance limit. By contrast, 1 m200 m wires replicate the sub-micron buried interconnects already deployed in high-density pixel detectors, biosensing micro-electrodes, and microwave striplines for NV-center control [9, 22, 23]. Studying both geometries therefore spans the full electrical landscape of laser-modified diamond; demonstrating a consistent correlation between Raman metrics and resistance across pads and wires is essential for future closed-loop fabrication. Accordingly, pads and wires of equal length were fabricated at varying laser scanning speeds to produce different degrees of graphitization, and thus a controlled range of electrical conductivities.
@@
Figures 1b and c show optical micrographs along with photoluminescence (PL) and Raman maps (sp3 and sp2) for two laser-written graphitic wires fabricated at scan speeds of 20 and 200 m s-1. Of note, these maps were acquired post-fabrication, as complete datasets including the pristine diamond were required for the study. The wire written at 20 m s-1 exhibits a resistance of , whereas the 200 m s-1 wire shows , corresponding to an difference (see Figure 4b for complete set of conductivity measurements). Details of the graphitic pads are provided in Supplementary Note 1. Under optical illumination, the laser-written track appears dark relative to the surrounding single-crystal diamond. This contrast arises because conversion from transparent, wide-band-gap sp3 diamond to disordered sp2 carbon markedly increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25]. Therefore, the degree of darkening can be used as a first qualitative indication of graphitization. However, it conflates absorption, scattering, debris, and illumination artifacts, and it is not phase-specific. For this reason, PL and hyperspectral Raman mapping are regarded as more reliable approaches, since they provide phase-specific, quantitative, and spatially resolved indicators of graphitization.
In the PL maps (Figure 1b,c), the laser-written graphitized regions appear dark because the processing introduces non-radiative recombination pathways that quench the broadband fluorescence. While tracking broadband emission can provide valuable information, monitoring the depletion of the sp3 Raman peak yields higher spatial resolution and stronger phase-specific contrast between pristine diamond and graphitized areas. Indeed, local graphitization converts sp3–bonded diamond into sp2 carbon, depleting the sp3 signal and producing the pronounced contrast. Of note, the sp3 maps were constructed by integrating the hyperspectral signal across the first-order diamond peak (1325-1340cm-1) [24, 26], while the sp2 maps were obtained by integrating over 1575-1610 cm-1 [24]. In the latter, since graphitization leads to the formation of sp2-bonded carbon domains it results in a marked increase in the G-band intensity.
1 Spectral characterization of Laser-written Electrodes
@@
As shown in Figure 1, one of the limitations of both brightfield microscopy and broadband PL imaging is that despite being good techniques for qualitative assessment of the fabrication they present limitations when used to provide quantitative information. As discussed in Supplementary Notes 3 and 4, PL could in principle provide quantitative evaluation of sp3 to sp2 conversion but, due to the broad integration, is very sensitive to other, unrelated phenomena, reducing quantitative fidelity. For this reason, we assessed the suitability of Raman spectroscopy for this task.
1.1 Single-band characterization
@@
To determine how the degree of graphitization affects the sp3 to sp2 transformation, we performed hyperspectral Raman mapping on every graphitic wire and pad fabricated at distinct scan speeds (see Materials and Methods). Spectra belonging to the written features and to the surrounding pristine diamond were separated pixel-by-pixel and averaged. Figures 2a and b display representative spectra for (a) the unmodified single-crystal diamond and (b) a fully graphitized pad fabricated at a laser scan speed of 100 m s-1. The pristine crystal is dominated by the sp3 line at 1332 cm-1. In contrast, the graphitic structure written with low laser scan speed exhibits depleted sp3 intensity and is governed by the disorder-induced D band (1350 cm-1) together with the sp2 band (graphitic G band (1580 cm-1)), evidencing almost complete conversion to sp2 carbon.
@@
Figure 2c compiles the speed-dependent averaged spectra of the laser-written graphitic wires. The spectral data normalized across all the samples to the sp3 peak (left panel) and the sp2 band (right panel), allowing the two regimes to be compared on equal footing across writing speeds. At the highest velocity tested (500 m s-1) the spectrum is sp3-dominated and the D and G bands are barely discernible, implying that only a small fraction of lattice sites have been converted to sp2. Progressively reducing the speed, and thereby increasing the deposited energy density per unit length, reverses this balance: the D and sp2 bands intensify while the sp3 line diminishes, signaling enhanced graphitization. The evolution is summarized quantitatively in Figure 2d, which plots the normalized integrated intensities and as well as their ratio versus laser scan speed. depletes with decreasing speed, whereas increases, except for a modest downturn below 50 m s-1, where excessive dose does not yields higher sp2 band intensity. The intensity ratio therefore spans almost a factor of three, from 0.9 at speeds 150 m s-1 to 0.3 at 500 m s-1, establishing laser scan speed as a single-parameter handle for tuning the local phase composition. A slight broadening of the sp3 line is also observed (see Supplementary Note 6), consistent with lattice strain introduced during partial sp2 formation.
1.2 Multi-band characterization
@@
Laser-written tracks in diamond contain both unconverted sp3-bonded lattice and newly formed sp2-bonded carbon. Within a confocal voxel, these phases typically coexist as sub-micron domains, so a single Raman spectrum is an average over a heterogeneous local volume. In the previous section we showed that the degree of graphitization can be inferred from the intensity of the sp3 line. However, relying on a single peak can be fragile in practice because the apparent intensity can be influenced by various factors such as laser power fluctuations, defocusing, local topography and debris, baseline fluorescence, or strain-induced shifts/broadening.
@@
To mitigate these effects, ratio metrics that combine several informative bands are usually preferred over single-peak readouts. In our case the most natural choice is to use the sp2/sp3 ratio, containing the information carried by the G/D complex and the 1332 cm-1 diamond line. Figure 3 illustrates this for a representative wire written at 50 m s-1: a 2D map of the ratio provides robust, self-normalized contrast across the field, enabling empirical thresholds that segment the scene into non-graphitized, partially graphitized, and fully graphitized classes. Class-averaged spectra make the assignments transparent: pristine regions show a narrow sp3 line accompanied by NV0 fluorescence, partially graphitized pixels exhibit concurrent sp3 attenuation and the emergence of D/G features, and fully graphitized regions are dominated by broad D/G bands with a strongly depleted 1332 cm-1 line.
@@
While ratio maps remove much of the acquisition variability, they still require manual selection of specific bands. Hyperspectral unmixing offers a label-free alternative that exploits the entire spectrum and returns per-pixel composition in the form of abundances. As described in Materials and Methods, endmembers are extracted directly from the data using well established unmixing methods, and abundances are then estimated per pixel, which enables a direct fractional interpretation without the need for external references.
@@
Applying this pipeline over a broad region of interest (1000–1800cm-1) yields endmembers that, while not identical to the actual reference spectra presented in Figure 2), preserve the expected relevant features: the diamond-associated endmember is characterised by a sharp 1332 cm-1 line and the NV0 contribution in the window, whereas the graphite-associated endmember exhibits a broad D/G complex with a suppressed diamond line. Importantly, this decomposition is entirely computational and label-free. The resulting abundance maps closely track the simpler sp3-window intensity maps, which could be anticipated given that the 1332 cm-1 line is the most intense and spatially specific feature, but provide a unified, quantitative representation across the whole spectrum.
@@
To test sensitivity to the dominant diamond line, we repeated the analysis after excluding 1332 cm-1 (spectral ROI = 1345–1800 cm-1). In this case, the diamond-associated endmember is defined primarily by the NV0 contribution within the spectral ROI, while the complementary endmember remains the broad G/D complex. The corresponding abundance maps retain the expected spatial structure but exhibit lower signal-to-noise, as anticipated when the most intense, phase-specific line is removed. This cropped-ROI approach is nevertheless useful in scenarios where the diamond line saturates the detector, where strong local strain or temperature gradients shift or distort the 1332 cm-1 feature beyond a fixed window, or when one wishes to decouple the analysis from sp3-dominated contrast (e.g. to emphasize changes within the sp2 manifold).
@@
Overall, the unmixing results reinforce the interpretation from ratio maps while removing the need to hand-pick bands: knowing only that two materials are present, the algorithm discovers representative endmembers and delivers per-pixel abundance maps that agree with the PL/Raman contrasts. These abundance maps offer a compact, quantitative metric for automated thresholding and closed-loop control (e.g. stop conditions based on sp3 depletion or sp2 abundance). Yet, in the particular case of monitoring graphitization in diamond, where the bands of interest are few and well known, hyperspectral unmixing does not provide a significant performance advantage over band-specific metrics.
2 Raman Monitoring of Electrical Performance

### v2

@@ Abstract
Femtosecond laser fabrication enables the creation of a wide range of devices, but its scalability and yield can be limited by the lack of real-time, in-situ monitoring tools. In particular, there is a strong need for metrics that directly correlate with device performance. Raman microspectroscopy provides a non-destructive route for in-situ characterization. Here, we demonstrate its potential to assess the electrical performance of laser-written graphitic electrodes in diamond. By combining hyperspectral mapping with electrical testing, we show that depletion of the 1332 cm-1 sp3 Raman line serves as a monotonic and robust predictor of resistance, offering clear advantages over commonly used spectral features. We further introduce hyperspectral unmixing as a label-free approach to identify relevant spectral signatures in fabrication processes where Raman markers are less defined. Importantly, the methodology we present is not restricted to diamond but can be adapted to other host materials and functionalities, offering a practical path toward specification-driven fs-laser microfabrication.
1 Department of Engineering Science, University of Oxford, Parks Road, Oxford OX1 3PJ, UK
2 Department of Materials, University of Oxford, Parks Road, Oxford OX1 3PH, UK
∗Corresponding authors: xingrui.cheng@eng.ox.ac.uk,
alvaro.fernandezgaliana@eng.ox.ac.uk
1 Introduction
Femtosecond laser writing enables direct, three-dimensional patterning of embedded features in transparent hosts via nonlinear absorption—multi-phonon ionization and Zener breakdown—that initiates localized structural transformation [gattass2008femtosecond, musgraves2011laser]. The technique has found particular application in processing of diamond [britel2025comparative, Jeschke1999microscopic, Ashikkalieva2016, Kononenko2016, sotillo2016], permitting fabrication of buried and surface carbon wires and electrode architectures with arbitrary geometry for electrically conductive devices [Sun2014], such as radiation detectors [komlenok2016diamond, Lagomarsino2013]. Yet, one of the main current limitations for scaling manufacturing of these laser-written devices is the lack of feedback mechanisms that can be integrated for real-time, in-situ process monitoring and control. Indeed, most current structural analyses remain destructive or ex-situ (e.g.,TEM [Ashikkalieva2019, Salter2024], SEM [Ashikkalieva2016], and X-ray diffraction microscopy [Yin2024]).
Raman microspectroscopy, based on inelastic light scattering, has been extensively used for non-destructive identification, structural characterization, and monitoring of chemical and physical properties across materials science, chemistry, and biology [fernandez2024fundamentals]. In recent years, significant efforts have also focused on developing advanced computational approaches for spectral analysis, further enhancing and broadening its range of applications.
Given its ability to differentiate the spectral signatures of sp3 versus sp2 bonding, crystallinity, and local strain environment, Raman spectroscopy is ideal for non-invasive, in-situ, spatially resolved characterization of carbon systems [khomich2022raman, Sotillo2018]. In carbon systems, Raman signatures provide direct markers of bonding and structural order: graphitic sp2 domains give rise to the G band near , while disorder or finite crystallite size activates the D band at [Tuinstra1970, Ferrari2000]. In contrast, single-crystal diamond exhibits a sharp first-order mode at [Knight1989, Ferrari2004]. Throughout this work, we denote these features as the sp3 line (), sp2 band (), and D band (), providing a consistent basis for assessing carbon phase composition, crystallinity, and defect populations in laser-written structures.
@@ 1 Introduction
Here, we use Raman microspectroscopy to systematically assess fs-laser-written graphitic electrodes in diamond, and compare its performance with brightfield and photoluminescence imaging. Leveraging phase-specific vibrational modes of carbon, we perform direct mapping of surface graphitization through explicit band assignment and label-free spectral deconvolution, and correlate the spectral signatures to the device performance (i.e., electrical conductivity). Beyond offering detailed insights into graphitization monitoring in diamond, the methodology outlined here is broadly transferable: it can be applied to evaluate Raman spectroscopy as a performance probe in diverse device platforms, since none of its principles are inherently diamond-specific.
2 Results and Discussion
@@
To model the principal conductive elements in diamond devices, we use a femtosecond laser to fabricate graphitic pads and wires on the surface of a CVD diamond sample, as shown in Figure 1a. Pads wider than 25 m and 200 m long reproduce the large‐area contact zones required for wire bonding or subsequent metallization, where sheet resistance and adhesion set the performance limit. By contrast, 1 m 200 m wires replicate the sub-micron buried interconnects already deployed in high-density pixel detectors, biosensing micro-electrodes, and microwave striplines for NV-center control [Salter2024, Bloomer2020, FORNERIS2017]. Studying both geometries spans the full electrical landscape of laser-modified diamond. Moreover, pads and wires of equal length were fabricated at varying laser scanning speeds to produce different degrees of graphitization, and thus a controlled range of electrical conductivities. For each of these structures, brightfield transmission microscopy images were captured, along with photoluminescence (PL) and backscattered Raman, collected using our integrated confocal microscope (see Materials and Methods).
@@
Figures 1b and c show these optical micrographs along with PL and Raman maps (sp3 and sp2) for two laser-written graphitic wires fabricated at scan speeds of 20 and 200 m s-1. The wire written at 20 m s-1 exhibits a resistance of , whereas the 200 m s-1 wire shows , corresponding to an difference (see Figure 4b for complete set of conductivity measurements). Details of the graphitic pads are provided in Supporting Note 1. Under optical illumination, the laser-written track appears dark relative to the surrounding single-crystal diamond. This contrast arises because conversion from transparent, wide-band-gap sp3 diamond to disordered sp2 carbon markedly increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [prelas1997handbook, Wang2000]. Therefore, the degree of darkening can be used as a first qualitative indication of graphitization. However, it conflates absorption, scattering, debris, and illumination artifacts, and it is not phase-specific. For this reason, PL and hyperspectral Raman mapping are regarded as more reliable approaches, since they provide phase-specific, quantitative, and spatially resolved indicators of graphitization.
In the PL maps (Figure 1b,c), the laser-written graphitized regions appear dark because the processing introduces non-radiative recombination pathways that quench the broadband fluorescence. Yet, it is not phase-specific and can be easily affected by broadband emission, as shown in Figure 1c and detailed in Supporting Note 2. On the other hand, Raman mapping provides phase-specific contrast. Local graphitization converts sp3 diamond to sp2 carbon, depleting the sp3 line and enhancing the G band; accordingly, we construct sp3 maps by integrating 1325–1340 cm-1 and sp2 maps by integrating 1575–1610 cm-1 [prelas1997handbook, Kononenko2008]. These phase-specific features (sp3 line, sp2 band, D band) enable quantitative, cross-structure comparison of phase content, crystallinity, and defect populations in laser-written carbon [Sotillo2018, Kononenko2008, Ashikkalieva2019]. Note that, as shown in (Figure 1c), Raman signal can also be contaminated by non-specific fluorescent background, an effect that can oftentimes be minimized with baseline correction and other techniques [wei2015review].
2.1 Spectral characterization of Laser-written Electrodes
@@
As shown in Figure 1, one of the limitations of both brightfield microscopy and broadband PL is that despite being good techniques for qualitative assessment of the fabrication they present limitations when used to provide phase-specific quantitative information. In the case of PL, it can in principle be used for quantitative sp3-to-sp2 conversion evaluation, but its sensitivity to broadband emission not stemming from the process of interest reduces its fidelity. This is further discussed in Supporting Notes 3 and 4. Therefore, we assessed the suitability of Raman spectroscopy for this task.
In graphitic carbons, the in-plane first-order phonon (“G band”) appears near , while disorder or finite crystallite size activates the defect-induced “D band” around [Tuinstra1970, Ferrari2000]. In single-crystal diamond, the zone-center optical phonon of symmetry gives a sharp first-order line at [Knight1989, Ferrari2004]. These emphasize the phase origin of each peak and facilitates quantitative comparison across laser-written structures, enabling identification and quantification of carbon phases, crystallinity, and defect populations [Sotillo2018, Kononenko2008, Ashikkalieva2019].
2.1.1 Single-band characterization
@@
Using the integrated confocal microscope, we collected hyperspectral Raman maps on every graphitic wire and pad fabricated at distinct scan speeds (see Materials and Methods). Figures 2a and b display representative spectra for (a) the unmodified single-crystal diamond and (b) a fully graphitized pad fabricated at a laser scan speed of 100 m s-1. The pristine crystal is dominated by the sp3 line at 1332 cm-1. In contrast, the graphitic structure written with low laser scan speed exhibits depleted sp3 intensity and is governed by the disorder-induced D band (1350 cm-1) together with the sp2 band (graphitic G band (1580 cm-1)), evidencing almost complete conversion to sp2 carbon.
@@
Figure 2c compiles the speed-dependent averaged spectra of the laser-written graphitic wires. The spectral data normalized across all the samples to the sp3 peak (left panel) and the sp2 band (right panel), allowing the two regimes to be compared on equal footing across writing speeds. At the highest velocity tested (500 m s-1) the spectrum is sp3-dominated and the D and G bands are barely discernible, implying that only a small fraction of lattice sites have been converted to sp2. Progressively reducing the speed, and thereby increasing the deposited energy density per unit length, reverses this balance: the D and sp2 bands intensify while the sp3 line diminishes, signaling enhanced graphitization. The evolution is summarized quantitatively in Figure 2d, which plots the normalized integrated intensities and as well as their ratio versus laser scan speed. depletes with decreasing speed, whereas increases, except for a modest downturn below 50 m s-1, where excessive dose does not yield higher sp2 band intensity.
2.1.2 Multi-band characterization
@@
Laser-written electrodes in diamond contain both unconverted sp3-bonded lattice and newly formed sp2-bonded carbon. Within a confocal voxel, these phases typically coexist as sub-micron domains, so a single Raman spectrum is an average over a heterogeneous local volume.
In the previous section we showed that the degree of graphitization can be inferred from the intensity of certain bands (e.g., sp3 line). However, relying on a single peak can be fragile in practice because the apparent intensity can be influenced by various factors such as laser power fluctuations, defocusing, local topography and debris, baseline fluorescence, or strain-induced shifts/broadening.
@@
To mitigate these effects, ratio metrics that combine several informative bands are usually preferred over single-peak readouts. In our case the most natural choice is to use the sp2/sp3 ratio, containing the information carried by the G/D complex and the 1332 cm-1 diamond line. Figure 3 illustrates this for a representative wire written at 50 m s-1.
2.1.3 Hyperspectral unmixing
@@
While ratio maps remove much of the acquisition variability, they still require manual selection of specific bands. Therefore, we also introduce the use of hyperspectral unmixing as a label-free alternative for fabrication processes where the bands of each phase might not be known. As described in Materials and Methods, we model each spectrum as a linear mixture of endmembers and use well-established linear unmixing methods, such as vector component analysis (VCA) [nascimento2005vertex] to estimate these endmembers directly from the data. Then, abundances are estimated per pixel solving the linear inverse problem. In particular, for the results shown in Figure 3, we impose endmembers and enforce both non-negativity, and the sum-to-one of the abundances at every pixel using fully constrained least squares (FCLS) [heinz2001fully], which enables a direct fractional interpretation without the need for external references.
@@
As shown in Figure 3e, when applied over a broad region of interest (1000–1800cm-1) unmixing yields endmembers that, while not identical to the actual reference spectra presented in Figure 3), preserve the expected relevant features: the diamond-associated endmember is characterised by a sharp 1332 cm-1 line, whereas the graphite-associated endmember exhibits a broad D/G complex with a suppressed diamond line. Additionally, the diamond-associated endmember also features a broader band centered around cm-1, which is consistent with the zero-phonon line of neutral single nitrogen-vacancy defects (NV0) in diamond [Manson2013]. The resulting abundance maps also closely track the simpler sp3-window intensity maps (Figure 3b).
@@
To test sensitivity to the dominant diamond line, we repeated the analysis after excluding 1332 cm-1 (spectral ROI = 1345–1800 cm-1). In this case, the diamond-associated endmember is defined primarily by the NV0 contribution within the spectral ROI, while the complementary endmember remains the broad G/D complex. The corresponding abundance maps retain the expected spatial structure but exhibit lower signal-to-noise (SNR), as anticipated when the most intense, phase-specific line is removed.
@@
Overall, the unmixing results reinforce the interpretation from ratio maps while removing the need to hand-pick bands: knowing only that two materials are present, the algorithm discovers representative endmembers and delivers per-pixel abundance maps that agree with the PL/Raman contrasts.
2.2 Raman Monitoring of Electrical Performance